# Eigen-Value: Efficient Domain-Robust Data Valuation via Eigenvalue-Based Approach

### v2

@@ 5 Experiments
We evaluate EV in three parts. (1) Cross-domain data removal and point addition. We compute values on a source domain and measure performance on a different target domain to test whether the valuation is useful for selection. (2) Stability and efficiency. To assess stability, we repeatedly alter a small subset of training samples and measure the variance in the resulting value rankings. We also compare computation time across methods. (3) Qualitative analysis. Top-ranked samples capture semantically invariant features of each class as well as exhibit broader dispersion in the embedding space, as confirmed in the qualitative top-3 examples. In contrast, low-ranked samples form redundant clusters and often miss such robust cues, explaining why EV enhances generalization under domain shift. All valuations are computed using our implementation based on the OpenDataVal Jiang et al. [2023]. Our code is available at https://github.com/MLAI-Yonsei/Eigen-Value.
@@ 5.1 Experimental Settings
We compare our proposed method against several baseline approaches for data valuation: (a) Random: Assigns data values randomly from a uniform distribution $U(0,1)$. (b) InfluenceFunction Feldman and Zhang [2020]: Estimates the influence of an individual training example on the validation dataset by computing closely related sub-sampled influence. (c) Deviation: Compute data values using the distributionally robust generalization error (DRGE) based on NTK. (d) LAVA: A model-agnostic data valuation method that utilizes the class-wise Wasserstein distance. (e) KNN Shapley: Computes Shapley values using the K-Nearest Neighbors (KNN) approach. (f) Data-OOB: Employs the Out-of-Bag (OOB) technique to estimate data values. (g) Eigen-Value: Our proposed approach EV, we conducted experiments applying other methods (LAVA, KNN Shapley, Data-OOB).
Datasets.
We conduct experiments on the following real-world datasets: (a) CIFAR-10 Krizhevsky et al. [2009]: A widely used image classification dataset consisting of 60K images across 10 classes. (b) CIFAR-10 C Hendrycks and Dietterich [2019]: A variant of CIFAR-10 where common corruptions are applied, resulting in an image dataset with a distribution shift from the original data. (c) VLCS Fang et al. [2013]: A dataset comprising images from four distinct domains (VOC2007, LabelMe, Caltech101, SUN09), all sharing the same label space. (d) Amazon Reviews Hou et al. [2024]: Amazon user product review data, which is organized into domains by product category. We convert the original 5-point ratings into three sentiment classes. In this work, we focus on the Books, Electronics, and Home and Kitchen domains. (e) ImageNet Deng et al. [2009]: A large-scale dataset of labeled natural images spanning thousands of object categories, widely used for visual recognition research. In this work, we use several of its derived domains: V2 Recht et al. [2019], S Gao et al. [2022], R Hendrycks et al. [2021a], and A Hendrycks et al. [2021b]. (f) DomainNet Peng et al. [2019]: A benchmark dataset designed to evaluate cross-domain generalization, comprising images from six distinct domains covering the same set of object categories.
Setting.
Our method views the upper bound of OOD loss as a utility to pick ID samples that best improve cross-domain generalization. We computed EV scores from normalized ID embeddings, combined them with ID loss-based valuations. Based on the integrated scores, we curated a subset of the ID data to train a logistic regression model evaluated on a target domain. For example, in VLCS, we designate SUN09 as the target domain. Data valuation and validation use the remaining three domains, and testing is performed solely on SUN09. Throughout this process, we use embeddings extracted from either ResNet50 He et al. [2016] and ViT-B/16 Dosovitskiy et al. [2020] for image-based datasets, and RoBERTa-base Liu et al. [2019] for text-based datasets such as Amazon Reviews.
@@ 5.2 Cross Domain Experiment
In the data removal experiment, we evaluate whether a valuation method can correctly identify low utility samples. From a pool of 2K ID data points, we randomly sample 1K for training and compute the value of each sample using each method. For every valuation baseline, we discard 50% of samples assigned the highest values and train the model using only the remaining half. The model is then evaluated on OOD data. Since the retained training set is composed of data deemed less valuable, a larger accuracy drop indicates that the method was more effective at flagging uninformative samples. As shown in Table 1, augmenting each baseline with EV consistently reduces the performance drop. Notably, EV + KNN Shapley achieves the best performance in all but two domains, where EV + Data-OOB outperforms all other approaches. These results highlight that EV provides a clearer guarantee of OOD robustness than existing alternatives such as Deviation.
To examine scalability beyond controlled small-scale settings and to test performance on more realistic and complex distributions, we further apply the same protocol to two large-scale benchmarks, ImageNet and DomainNet. Specifically, we subsample 30K images from ImageNet and 10K from DomainNet, compute per-sample values with all baselines in Table 1, and repeat the data removal procedure. For scalability reasons, we omit LAVA and Deviation due to their prohibitive computational costs. On ImageNet, we evaluate across four domain shifts (V2, Sketch, Rendition, Real), while on DomainNet we report results averaged across six shifts (Clipart, Infograph, Painting, Quickdraw, Real, Sketch), with per-domain details provided in Appendix C.2. As summarized in Table 2, the EV-augmented methods once again yield the lowest error in almost every target domain, confirming that our approach scales effectively and delivers superior robustness under substantial distribution shifts.
@@
In the point addition experiment, we simulate a scenario where additional ID data is incrementally incorporated into training to assess its effect on OOD generalization. We begin by randomly sampling 2K ID data points and computing data values with each valuation method. From this pool, we construct an initial training set of 1K samples and then gradually expand it by adding the highest-valued samples from the remaining 1K points in descending order, retraining the model after each addition. This process is repeated on CIFAR-10 (evaluated on CIFAR-10 C), VLCS, and Amazon Reviews, thereby covering both vision and text domains. Figure 4 reports the performance after each addition step. Solid bars represent the baseline methods, and hatched bars indicate their results when EV is applied. Across domains, EV consistently yields higher accuracy and robustness to distribution shifts. These results show that incorporating EV into data valuation effectively guides data selection toward more informative and resilient samples. Overall, the point-addition experiment demonstrates that EV provides a principled criterion for selecting additional ID data that yields the greatest benefit under OOD evaluation. This makes EV particularly suitable for continual or online learning scenarios, where deciding which incoming samples to prioritize is a critical challenge.
Figure 5: Stability under training-set perturbations. We conduct a valuation on 300 CIFAR-10 samples five times, keeping 290 fixed and resampling 10. Deviation’s data-value ranking exhibits a standard deviation comparable to random selection, whereas EV yields stable, efficient rankings while retaining ID-based valuation and strong OOD performance.
@@ 5.3 Stability and Efficiency of Eigen-Value
A good data valuation method must be stable to small changes in the subset. Deviation estimates worst-case distribution error using NTK analysis by constructing a separate leave-one-out NTK matrix for each sample, which amplifies sensitivity to dataset composition and leads to unstable rankings. To evaluate stability, we fixed 290 out of 300 training samples and randomly replaced the remaining 10, repeating this procedure five times to compute the standard deviation of rankings. Ideally, if only a small portion of the dataset changes, the rankings of the fixed samples should vary within that range. However, as shown in Figure 5, Deviation exhibits much larger fluctuations, with the standard deviation of rankings approaching that of random selection. EV quantifies each sample’s contribution to domain discrepancy through covariance eigenvalues and produces consistent rankings. This stability makes EV more reliable for real-world data markets.
(a) Top-3 samples selected by Data-OOB. Some images fail to capture invariant structures (e.g., dogs without sleds or unclear pulling).
(b) Top-3 samples selected by EV + Data-OOB. EV consistently highlights dogs visibly pulling a sled.
Figure 7: Qualitative comparison of the top-3 ranked images in the dog sled class, selected according to data values from (a) Data-OOB and (b) EV + Data-OOB.
Figure 8: PCA projection of top-1K and bottom-1K CIFAR-10 samples selected by Data-OOB valuation with and without EV. Incorporating EV leads to a greater variance gap (Top: 1.09, Bottom: 0.67 in EV + Data-OOB vs. Top: 0.38, Bottom: 0.60 in Data-$O$), making it easier to identify widely dispersed, high-variance samples. Such samples better cover diverse features, which improves generalization to OOD data with different distributions.
@@
Another novelty of our approach is that it performs data valuation operations, yielding high performance within a short time. Although Deviation is a data valuation method that considers OOD data, it inverts an $n-1\times n-1$ matrix for each data point, resulting in a prohibitive computational complexity. This excessive computational burden, which grows cubically with dataset size, makes the method practically infeasible and limits its scalability in real-world applications. In contrast, our method adds only a small amount of additional computation compared to existing methods while demonstrating superior OOD performance. As shown in Figure 6, the methods augmented with EV achieve better OOD performance with minimal overhead of computing approximate eigenvalues for 2K samples, and take less than 1 second. In contrast, Deviation requires nearly 30 minutes and lacks OOD robustness.
@@ 5.4 Qualitative Analysis
In our PCA projection analysis, we visualize the top and bottom 1K samples of ImageNet as ranked by each valuation method, using the valuation results reported in Table 2. Figure 8 shows these samples, selected by Data-OOB and EV + Data-OOB, projected onto the top three principal components. For robust OOD performance, it is preferable to train on samples that are broadly distributed in the feature space, rather than narrowly clustered. To assess the impact of EV, we compared the variance of high- and low-value samples. Incorporating the EV term increased the variance of top-ranked as well as enlarged the gap between top- and bottom-ranked groups (Top: 1.09 vs. Bottom: 0.67 in EV + Data-OOB, compared to Top: 0.38 vs. Bottom: 0.60 in Data-OOB). In fact, Data-OOB even assigns lower variance to top-ranked samples than to bottom-ranked ones, indicating that it sometimes fails to prioritize diverse and informative examples. By contrast, EV consistently highlights widely dispersed, high-variance samples, while keeping low-value samples more concentrated. As a result, models trained on EV-selected data are exposed to richer and more representative features, which enhances their robustness under distribution shifts.
@@ 6 Conclusion
In this paper, we propose Eigen-Value (EV), an efficient data valuation framework for OOD robustness. By approximating domain discrepancy via eigenvalues and perturbation theory, EV estimates the marginal contribution of each sample to OOD loss. Integrated with existing ID-based methods, it enables OOD-aware data selection without requiring any OOD data. Comprehensive cross-domain experiments on vision and text datasets demonstrate that EV consistently enhances domain robustness while maintaining stability and low computational cost. Qualitative analyses further reveal why EV improves robustness, showing that it prioritizes diverse, invariant features that are critical under distribution shifts. By shifting the focus from model- to data-centric OOD robustness, EV offers a scalable solution with theoretical guarantees linking spectral properties to OOD generalization. Together, these results establish EV as a practical and reliable tool for real-world applications, where robust and efficient data valuation is essential.
Appendix A Theorerical Analysis
@@ A.1 Estimating Domain Discrepancy using Eigenvalue Shifts Induced by NCE
Thus, the ratio can be approximated by Hessiansian of each distribution ($H_{\text{OOD}},H_{\text{ID}}$)
@@ A.2 Using Logistic Regression Hessian as a Covariance Approximation
Thus, the Hessian of the NCE (a variant of logistic regression) is upper boundedthe by covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$
@@ A.3 Approximating Marginal Contributions of the Eigenvalue Term
$\delta_{\max}^{(k)}:=u_{\max}^{\top}\Delta_{k}\,u_{\max},\quad\delta_{\min}^{(k)}:=u_{\min}^{\top}\Delta_{k}\,u_{\min}.$
@@
$A=\lambda_{\max}(\Sigma_{\textnormal{ID}})\times\sqrt{d}+\sqrt{d^{2}-d},\quad B=\lambda_{\min}(\Sigma_{\textnormal{ID}}).$
@@
$f(\Sigma_{-k})-f(\Sigma_{\textnormal{ID}})\approx\frac{\sqrt{d}\times\delta_{\max}^{(k)}}{B}-\frac{A\times\delta_{\min}^{(k)}}{B^{2}}=\frac{\sqrt{d}\times\delta_{\max}^{(k)}}{\lambda_{\min}(\Sigma_{\textnormal{ID}})}\\ -\frac{(\lambda_{\max}(\Sigma_{\textnormal{ID}})\times\sqrt{d}+\sqrt{d^{2}-d}\,)\times\delta_{\min}^{(k)}}{\lambda_{\min}(\Sigma_{\textnormal{ID}})^{2}}.$ (21)
@@
Conclusion: The derivations above demonstrate how domain discrepancy can be bounded by the eigenvalue ratio of the Hessians, how the OOD covariance matrix can be related to the ID covariance matrix, and how perturbation analysis yields an approximation of the marginal data value.
Appendix B Additional Experiment Setting

### v3

@@ 5 Experiments
We evaluate EV in three parts. (1) Cross-domain data removal and point addition. We compute values on a source domain and measure performance on a different target domain to test whether the valuation is useful for selection. (2) Stability and efficiency. To assess stability, we repeatedly alter a small subset of training samples and measure the variance in the resulting value rankings. We also compare computation time across methods. (3) Qualitative analysis. Top-ranked samples capture semantically stable class features and show wider dispersion in the embedding space, as seen in the qualitative top-3 examples. In contrast, low-ranked samples cluster redundantly and often lack such cues, explaining EV’s improved generalization under domain shift. All valuations are computed using our implementation based on the OpenDataVal [20]. Our code is available at https://github.com/MLAI-Yonsei/Eigen-Value.
@@ 5.1 Experimental Settings
We compare our method, EV, with several baseline approaches for data valuation: (a) Random: Assigns data values randomly from a uniform distribution $U(0,1)$. (b) InfluenceFunction [6]: Estimates the influence of an individual training example on the validation dataset by computing closely related sub-sampled influence. (c) Deviation: Compute data values using the distributionally robust generalization error (DRGE) based on NTK. (d) LAVA: A model-agnostic data valuation method that utilizes the class-wise Wasserstein distance. (e) KNN Shapley: Computes Shapley values using the K-Nearest Neighbors (KNN) approach. (f) Data-OOB: Employs the Out-of-Bag (OOB) technique to estimate data values. (g) Eigen-Value: Our proposed approach, EV, we conducted experiments applying other methods (LAVA, KNN Shapley, Data-OOB).
@@
Datasets.
We conduct experiments on the following diverse datasets: (a) CIFAR-10 [24]: A widely used image classification dataset containing 60k images across 10 classes. (b) CIFAR-10 C [13]: A variant of CIFAR-10 where common corruptions are applied, resulting in an image dataset with a distribution shift from the original data. (c) VLCS [5]: A dataset comprising images from four distinct domains (VOC2007, LabelMe, Caltech101, SUN09), all sharing the same label space. (d) Amazon Reviews [16]: Amazon user product review data, which is organized into domains by product category. We convert the original 5-point ratings into three sentiment classes. In this work, we focus on the Books, Electronics, and Home and Kitchen domains. (e) ImageNet [3]: A large-scale dataset of labeled natural images spanning thousands of object categories, widely used for visual recognition research. In this work, we use several of its derived domains: ImageNet-V2 (V2) [41], ImageNet-Sketch (S) [53], ImageNet-R (R) [12], and ImageNet-A (A) [14]. (f) DomainNet [39]: A benchmark dataset designed to evaluate cross-domain generalization, comprising images from six distinct domains covering the same set of object categories.
Setting.
Our method views the upper bound of OOD loss as a utility to pick ID samples that best improve cross-domain generalization. We computed EV scores from normalized ID embeddings, combined them with ID loss-based valuations. Based on the integrated scores, we curated a subset of the ID data to train a logistic regression model, which we optimized for 10 epochs using Adam with a learning rate of 0.001 and evaluated on the target domain. For example, in VLCS, we designate SUN09 as the target domain. Data valuation and validation use the remaining three domains, and testing is performed solely on SUN09. Throughout this process, we use embeddings extracted from either ResNet50 [11] and ViT-B/16 [4] for image-based datasets, and RoBERTa-base [31] for text-based datasets such as Amazon Reviews.
@@ 5.2 Cross Domain Experiment
In the data removal experiment, we evaluate whether a valuation method can correctly identify low utility samples. From a pool of 2k ID data points, we randomly sample 1k for training and compute the value of each sample using each method. For every valuation baseline, we discard 50% of samples assigned the highest values and train the model using only the remaining half. The model is then examined on OOD data. Since the retained training set is composed of data deemed less valuable, a larger accuracy drop indicates that the method was more effective at flagging uninformative samples. As shown in Table 1, augmenting each baseline with EV consistently reduces the performance drop. Notably, EV + KNN Shapley achieves the best performance in all but two domains, where EV + Data-OOB outperforms all other approaches. These results highlight that EV provides a clearer guarantee of OOD robustness than existing alternatives such as Deviation.
To examine scalability beyond controlled small-scale settings and to test performance on more realistic and complex distributions, we further apply the same protocol to two large-scale benchmarks, ImageNet and DomainNet. Specifically, we subsample 30k images from ImageNet and 10k from DomainNet, compute per-sample values with all baselines in Table 1, and repeat the data removal procedure. For scalability reasons, we omit LAVA and Deviation due to their prohibitive computational costs. On ImageNet, we evaluate across four domain shifts (V2, S, R, A), while on DomainNet we report results averaged across six shifts, with per-domain details provided in Appendix C.1. As summarized in Table 2, the EV-augmented methods once again yield the lowest error in almost every target domain, confirming that our approach scales effectively and delivers superior robustness under substantial distribution shifts.
@@
In the point addition experiment, we simulate a scenario where additional ID data are incrementally added to assess their effect on OOD generalization. We sample 2k ID data points, compute values with each method, and construct an initial 1k training set. The remaining 1k samples are then added in descending order of value, retraining the model after each step. This process is repeated on CIFAR-10 (tested on CIFAR-10 C), VLCS, and Amazon Reviews, covering both vision and text domains. Figure 3 shows performance after each addition, where solid bars denote baselines and hatched bars indicate their EV-augmented counterparts. Across domains, EV consistently achieves higher accuracy and stronger robustness to distribution shifts. These results confirm that EV effectively guides data selection toward informative and resilient samples, offering a principled criterion for continual or online learning, where prioritizing incoming data is critical.
Figure 4: Stability under training-set perturbations. We conduct a valuation on 200 CIFAR-10 samples five times, keeping 180 fixed and resampling 20. Deviation’s data-value ranking exhibits a standard deviation comparable to random selection, whereas EV yields stable, efficient rankings while retaining ID-based valuation and strong OOD performance.
@@ 5.3 Stability and Efficiency of Eigen-Value
A good data valuation method must be stable to small changes in the subset. Deviation estimates worst-case distribution error using NTK analysis by constructing a separate leave-one-out NTK matrix for each sample, which amplifies sensitivity to dataset composition and leads to unstable rankings. To evaluate stability, we fixed 180 out of 200 training samples and randomly replaced the remaining 20, repeating this procedure five times to compute the standard deviation of rankings. Ideally, if only a small portion of the dataset changes, the rankings of the fixed samples should vary within that range. However, as shown in Figure 4, Deviation exhibits much larger fluctuations, with the standard deviation of rankings approaching that of random selection. EV quantifies each sample’s contribution to domain discrepancy through covariance eigenvalues and produces consistent rankings. This stability makes EV more reliable for real-world data markets.
(a) Top-3 samples in the dog sled class selected by Data-OOB after data valuation. Some images fail to capture invariant structures (e.g., dogs without sleds or unclear pulling).
(b) Top-3 samples in the dog sled class selected by EV + Data-OOB after data valuation. EV consistently prioritizes images where dogs are clearly pulling a sled.
Figure 6: Qualitative comparison of the Top-3 ranked images in the dog sled class from ImageNet, selected according to data values from (a) Data-OOB and (b) EV + Data-OOB. For each method, we first compute data values of images and then extract the three highest-value samples within the dog sled class.
Figure 7: PCA projection of the top-1k and bottom-1k CIFAR-10 samples selected by Data-OOB with and without EV. Data-OOB alone produces an unexpected pattern where top samples have lower variance than bottom samples (0.38 vs. 0.60). With EV incorporated (EV + Data-OOB), the variance ordering becomes the opposite of what Data-OOB produces, with top samples showing larger dispersion than bottom samples (1.09 vs. 0.67). This helps the method identify broadly dispersed, high-variance samples that capture diverse features and improve OOD generalization.
@@
Another novelty of our approach is that it performs data valuation operations, yielding high performance within a short time. Although Deviation is a data valuation method that considers OOD data, it inverts an $n\times n$ matrix for each data point, resulting in a prohibitive computational complexity. This excessive computational burden, which grows cubically with dataset size, makes the method practically infeasible and limits its scalability in real-world applications. In contrast, our method adds only a small amount of additional computation compared to existing methods while demonstrating superior OOD performance. As shown in Figure 5, the methods augmented with EV achieve better OOD performance with minimal overhead of computing approximate eigenvalues for 2k samples, and take less than 1 second. In contrast, Deviation requires nearly 30 minutes and lacks OOD robustness.
@@ 5.4 Qualitative Analysis
In Figure 7, we analyze the top-1k and bottom-1k ImageNet samples ranked by Data-OOB, with and without EV, and project them onto the top three principal components. For robust OOD performance, it is preferable to train on samples that are broadly distributed rather than narrowly clustered. However, Data-OOB without EV shows an inverted variance pattern in which top-ranked samples are more tightly clustered than bottom ones (Top: 0.38 vs. Bottom: 0.60), indicating that it may overlook diverse, informative examples. Adding EV restores the expected structure: top samples exhibit substantially higher dispersion than bottom samples (Top: 1.09 vs. Bottom: 0.67). This shows that EV prioritizes broadly distributed, high-variance samples, leading to a more representative training subset and stronger robustness under distribution shifts.
@@ 6 Conclusion
In this paper, we propose Eigen-Value (EV), an efficient data valuation framework for OOD robustness. By approximating domain discrepancy via eigenvalues and perturbation theory, EV estimates each sample’s marginal contribution to OOD loss. Integrated with existing ID-based methods, it enables OOD-aware data selection without requiring additional OOD data. Extensive experiments on visual and textual datasets show that EV consistently enhances domain robustness while maintaining stability and low computational cost. Qualitative analyses further explain this improvement, showing that EV prioritizes diverse, invariant features critical under distribution shifts. By shifting the focus from model- to data-centric robustness, EV provides a scalable, theoretically grounded solution for real-world applications where efficient and reliable data valuation is essential.
Acknowledgment
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00457216).
Supplementary Material
A Theoretical Analysis
@@ A.1 Estimating Domain Discrepancy using Eigenvalue Shifts Induced by NCE
Thus, the ratio can be approximated by the Hessian of each distribution ($H_{\text{OOD}},H_{\text{ID}}$)
@@ A.2 Using Logistic Regression Hessian as a Covariance Approximation
Thus, the Hessian of the NCE (a variant of logistic regression) is upper bounded by the covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$
@@ A.3 Approximating Marginal Contributions of the Eigenvalue Term
$\delta_{\max}^{(k)}:=u_{\max}^{\top}\Delta_{k}\,u_{\max},\qquad\delta_{\min}^{(k)}:=u_{\min}^{\top}\Delta_{k}\,u_{\min}.$
@@
$A=\lambda_{\max}(\Sigma_{\textnormal{ID}})\times\sqrt{d}+\sqrt{d^{2}-d},\qquad B=\lambda_{\min}(\Sigma_{\textnormal{ID}}).$
@@
$\begin{split}&f(\Sigma_{-k})-f(\Sigma_{\textnormal{ID}})\approx\frac{\sqrt{d}\times\delta_{\max}^{(k)}}{B}-\frac{A\times\delta_{\min}^{(k)}}{B^{2}}\\[14.22636pt] &=\frac{\sqrt{d}\times\delta_{\max}^{(k)}}{\lambda_{\min}(\Sigma_{\textnormal{ID}})}-\frac{(\lambda_{\max}(\Sigma_{\textnormal{ID}})\times\sqrt{d}+\sqrt{d^{2}-d}\,)\times\delta_{\min}^{(k)}}{\lambda_{\min}(\Sigma_{\textnormal{ID}})^{2}}.\end{split}$ (21)
@@
The derivations above demonstrate how domain discrepancy can be bounded by the eigenvalue ratio of the Hessians, how the OOD covariance matrix can be related to the ID covariance matrix, and how perturbation analysis yields an approximation of the marginal data value. In addition, we examine the potential bias arising from dominant eigenvalues, as previously discussed in PCA-related analyses [43]. Our formulation indicates that robustness improves when the spectral distribution is balanced, with sufficiently large minimum eigenvalues and no excessively dominant maximum eigenvalues. This observation aligns with prior findings, suggesting that over-reliance on dominant directions can lead to biased representations, while a more uniform eigenvalue spectrum contributes to improved stability and generalization.
B Additional Experiment Setting